OpenStudy (anonymous):

if 300-0.002Q= 20+0.008Q and you have to make the Q the subject. does 300 change to -300 and does 20 also becomes negative|? i am strugling with the -+ changes

OpenStudy (anonymous):

300-0.002Q = 20+0.008Q -0.002Q-0.008Q = 20-300 -0.01Q = -280 negative sign cancels out on both side. Q = 280/ 0.01 Q = 28000